Garda purchased $670,000 of merchandise in August and expects to purchase $750,000 in September. Merchandise purchases are paid as follows: 15% in the month of purchase and 85% in the following month.

Compute cash payments for merchandise for September.
Answer #1

Cash payment more purchased in August=670000*85% =569500

Cash payment more purchased in september=750000*15% = 112500

Total cash payment for September =682,000